# Learning Heterogeneous Mixture of Scene Experts for Large-scale Neural Radiance Fields

### v1

@@ 4.2 Implement Details
In the MLP-based Switch-NeRF, the foreground network contains 8 NeRF experts configured with a Top-$1$ gating network and a capacity factor $C_{f}=1.0$, while the background is represented by a single NeRF network. Each NeRF expert is structured as a 7-layers MLP with 256 channels per layer. In training and inference, 256 coarse and 512 fine points are sampled per ray for the foreground network, whereas 128 coarse and 256 fine samples for the background network. The training utilizes 8 NVIDIA RTX 3090 GPUs, with a global batch size of 8192 rays in each iteration. The models are optimized by 500k iterations for each large-scale scene and then evaluated on validation images.
@@
In our Switch-NeRF++ framework, We utilizes multi-resolution hash encoding with 16 levels, each level with 2 feature channels. The maximum hash map size is $2^{19}$. For the Mega-NeRF and Matrixcity datasets, we use the same hash experts. For the larger Block-NeRF dataset and UrbanBIS dataset, we validate our design of heterogeneous mixture of hash experts (HMoHE) using both the same hash experts and heterogeneous hash experts. The hash gating network includes a multi-resolution hash encoding and a 3-layer MLP with 64 channels. We set 8 hash experts in our experiments for the foreground and one hash encoding for the background. The $\sigma$ prediction head is a 2-layer MLP and the color prediction head is a 3-layer MLP, both with 64 channels. We conduct our experiments on 4 NVIDIA RTX 3090 GPUs, with a batch size of 16384 rays per GPU. We use larger batch sizes to fully use the computation power of the GPUs since our network uses much less memory than MLP-based NeRF methods. We use coarse-to-fine sampling similar to Mega-NeRF, sampling 128 coarse samples and 128 fine samples along a ray in the foreground and half of them in the background.
@@
We use Adam optimizer to train Switch-NeRF and Switch-NeRF++ with a learning rate of $5\times 10^{-4}$ that exponentially decreases. The weight $\lambda$ of balance loss $\mathcal{L}_{b}$ is set as $5\times 10^{-4}$ for all experiments. We also use half-precision training and testing similar to Mega-NeRF.
@@
Baseline. We compare our method with the large-scale NeRF method Mega-NeRF [2]. We also adapt the Instant-NGP [5] into large-scale setting, named INGP. We use a foreground and a background network in the baseline. Each of them consists of one multi-resolution hash encoding and prediction heads. We set the hash map size as $2^{22}$ in the foreground network so it has the same network capacity as our method. The sampling strategy is the same as our Switch-NeRF++.
@@
TABLE I: The benchmark performances of our Switch-NeRF++ on the Mega-NeRF [2] dataset. In this table, the INGP is the Instant-NGP [5] trained with 100k steps. HMoHE* is trained with 50k steps and HMoHE is trained with 100k steps. Our method can achieve state-of-the-art results with only about a training time of 5.5h, which is an 8x acceleration compared to Switch-NeRF [4] and a 5x acceleration compared to Mega-NeRF [2].
@@ 4.3 Overall Benchmark Results
Mega-NeRF dataset. We evaluate the accuracy and efficiency of our framework on Mega-NeRF [2] dataset. The accuracy and training time are reported in Table I. The results of Mega-NeRF is quoted from their papers. In Table I, Mega-NeRF and Switch-NeRF are trained with about 30h and 44h respectively for 500k steps. Instant-NGP (INGP) [5] is trained with about 10h for 100k steps. Switch-NeRF++* is our method trained with 5.5h for 50k steps. Switch-NeRF++ is our method trained with 11h for 100k steps. The results show that our MLP-based Switch-NeRF can clearly outperform its counterpart Mega-NeRF, validating the effectiveness of our learnable decomposition and Mixture of Experts. Switch-NeRF++* trained with only 5.5 hours can already achieve state-of-the-art results on almost all the metrics. This is an acceleration of 8×8$\times$8 × compared to Switch-NeRF and 5×5$\times$5 × compared to Mega-NeRF, proving the superior efficiency and accuracy of our method. Our Switch-NeRF++ clearly outperforms INGP with less training time, showing the great scalability of our Switch-NeRF++ framework. The training time of our method is comparable to INGP with the same steps. One reason is that we implement an efficient dispatching method by integrating point dispatching into the hash operation. Another reason is that we set a similar number of parameters in INGP as our network. We provide qualitative results in Fig. 7. Our method can render images of a higher quality with cleaner and sharper details. The main results on the Mega-NeRF dataset clearly demonstrate that our Switch-NeRF++ can achieve better accuracy with superior efficiency comparing to Mega-NeRF, Switch-NeRF and INGP.
UrbanBIS dataset. We evaluate the accuracy and scalability of our method on the very large UrbanBIS [10] dataset. Table II demonstrate that our heterogeneous hash experts (Heter.) significantly outperform all the compared methods. Notably, our heterogeneous version shows substantial improvements over our homogeneous version (Homo.) across all metrics (PSNR, SSIM, and LPIPS), emphasizing the importance of considering scene heterogeneity when scaling to larger scenes. These experiments strongly support the scalability of our design of HMoHE in Switch-NeRF++ when applied to very large-scale scenes. Fig. 8(a) illustrates that our method produces images with enhanced sharpness and details.
@@
TABLE III: PNSR on MatrixCity [73] dataset. Switch-NeRF++ achieves an overall 1.75-point improvement in PSNR. The results of other methods are quoted from the MatrixCity paper. Please note that INGP also uses $2^{22}$ entries, whose number of parameters align with our network. These experiments demonstrate that our method can handle different types of scenes.
@@
4.4 Model Analysis
Effect of heterogeneous mixture of hash experts. We validate our heterogeneous hash design on Block-NeRF [3] and UrbanBIS [10] datasets. The scene in Block-NeRF dataset covers a $1km$ street. The scene in UrbanBIS dataset covers a $6.5km^{2}$ region. Both scenes are very large to evaluate the scalability of our heterogeneous design. The results in Table II clearly demonstrate the superior performance of our heterogeneous design compared to our homogeneous version (Homo.), particularly on UrbanBIS. Each of the experts in Switch-NeRF++ (Homo.) uses the default setting of the encodings in Instant-NGP [5]. In contrast, Switch-NeRF++ (Heter.) incorporates experts with varying resolution ranges, with minimum resolutions of the experts exponentially increasing from 16 to 512 and maximum resolutions exponentially increasing from 2048 to 16384. The ablation results strongly emphasize the significance of considering scene heterogeneity when scaling to larger scenes and provide clear evidence for the scalability of our method. Fig. 8(b) visualize the decomposition of the radiance field of UrbanBIS dataset of Switch-NeRF++ (Homo.). and Switch-NeRF++ (Heter.). These visualizations further validate that our heterogeneous design can effectively capture the different distributions of the large-scale scene.
@@ 4.4 Model Analysis
Gating network design in Switch-NeRF. We analyze different designs of the MLP-based gating network in Switch-NeRF [4]. Table VIII compares three configurations. The first is using a Linear gating, which processes the positional encoding PE(x) by only a trainable linear layer. The second is removing the LayerNorm in the gating network. The third is training the network without the gating loss $L_{b}$. As shown in the table, the Linear gating underperforms due to insufficient network capacity, validating that the MLP+Norm design in our gating network can substantially improves reconstruction accuracy. Notably, removing the gating loss $L_{b}$ prevents the training to converge, highlighting that $L_{b}$ is necessary for stabilizing the optimization of our sparse MoE system. This finding aligned with prior works  [11, 62].
Full dispatch in Switch-NeRF. We compare the inference accuracy of using uniform dispatch or using our full dispatch proposed in Sec. 3.5.1 on the trained Switch-NeRF. The Batch Prioritized Routing [63] and a capacity factor $C_{f}$ of $2.0$ or $4.0$ are used in the inference of the uniform dispatch. As shown in Table XII, uniform dispatch with $C_{f}=2.0$ suffers significant quality degradation. While increasing $C_{f}$ to $4.0$ partially improves the quality, it remains substantially inferior to full dispatch. Meanwhile, larger $C_{f}$ introducing extra computational overhead due to zero padding. Visual comparisons in Fig. 6 further highlight that the uniform dispatch renders images with much more artifacts, whereas our full dispatch produces complete and higher-quality outputs. This validates the effectiveness of using full dispatch in inference of Switch-NeRF.
Effect of the hash-based gating network. We replace the hash-based gating network in Switch-NeRF++ with an MLP gating network similar to that in Switch-NeRF [4] which consists of 4 linear layers with a width of 256. The input of the MLP gating network is the positional encoding [1]. We train them with 100k steps. As shown in Table IV, the accuracy drops by a large margin with the MLP gating network. This may be because the MLP suffers slow converging speed, and the representation power of the MLP gating network is inconsistent with the hash experts in Switch-NeRF. This experiment further confirms the importance of our design of the hash-based gating network.
TABLE VIII: Different gating network designs in Switch-NeRF on Sci-Art scene.
@@
TABLE IX: Comparison of Switch-NeRF and Mega-NeRF with different sub-networks on accuracy, test time, test memory and network parameters numbers. Switch-NeRF demonstrates better scalability than Mega-NeRF, increasing performance without a large increase of test memory and time.
@@
TABLE X: The ablation results of Switch-NeRF using different Top-$1$, Top-$2$ and capacity factors. A larger $C_{f}$ or employing Top-$2$ operation in the gating network in training can improve the accuracy slightly while consuming more training time and memory.
@@
TABLE XI: The ablation results comparing Switch-NeRF [4] and Mega-NeRF [2] on parameter number, time, memory and FLOPs. Although using slightly more training time and memory, Switch-NeRF consumes less memory in testing and has much less parameters while achieving a better accuracy.
[…]
Mega-NeRF 6935M 87.9s 5124M 30.7h 10.8M 0.79M
Switch-NeRF 5847M 110s 10182M 42.5h 4.53M 0.99M
TABLE XII: The testing time and accuracy of uniform dispatch with $C_{f}=2.0$ and $4.0$ and full dispatch. Using the full dispatch can achieves much better performance with less time usage.
@@
TABLE XIII: Impact of different number of gating operations on accuracy and efficiency in Switch-NeRF. Despite maintaining the same number of expert layers, Switch-NeRF with two gating operations fails to enhance rendering accuracy compared to the model with a single gating operation, while requiring increased memory usage and longer training and inference times.
@@
Effect of the number of experts. We analyze the effect of the number of NeRF experts in both Switch-NeRF [4]. Table IX reveals that even using 4 experts, Switch-NeRF already matches the accuracy of Mega-NeRF with 8 sub-networks. Scaling Switch-NeRF to 16 experts yields significant performance gains compared to 8 experts, while while incurring minimal additional computational and memory overhead during inference. This contrasts to the Mega-NeRF with 16 sub-networks, where Switch-NeRF not only scales much better and achieves a higher accuracy, but also reduces parameters and memory usage. The results demonstrate the superior scalability of Switch-NeRF in balancing model capacity and resource demands.
@@
Plane features without decomposition. The GP-NeRF [53] augments hash encodings with 2D ground plane features. It does not perform scene decomposition. The comparing results are shown in Table V. Switch-NeRF++ consistently outperform GP-NeRF, especially on the very large UrbanBIS [10] dataset. These experiments validate the importance of our learnable heterogeneous decomposition on large-scale scenes to better exploit network parameters.
@@
Vanilla Gaussian splatting. We compare with the vanilla Gaussian splatting method (3D-GS) [44] in Table VIIby applying 3D-GS on large-scale datasets. The results demonstrate that the vanilla 3D-GS struggles on large-scale scenes. It relies on a point cloud for initialization, while point clouds in large-scale scenes are typically noisy, contain outliers, and have significantly varying densities. The original 3D-GS has no special designs for these problems, resulting in poor initialization and training. Our method incorporates the learnable heterogeneous decomposition specially designed for large-scale scenes and can get much better results. Moreover, our method does not need any point clouds for initialization, making it more suitable for various large-scale reconstruction tasks.
Figure 9: The accuracy of our Switch-NeRF++ and other methods with respect to training steps. Our method continues to improve with more training steps.
TABLE XIV: Rendering time and memory. Our method speed up rendering by 16x compared to Switch-NeRF [4] and by 13x compared to Mega-NeRF [2] with much less memory.
Mega-NeRF Switch-NeRF Switch-NeRF++
Mem.$\downarrow$ 6935M 5847M 2857M
[…]

### v2

@@ 4.2 Implementation Details
In the MLP-based Switch-NeRF, the foreground network contains 8 NeRF experts configured with a Top-$1$ gating network and a capacity factor $C_{f}=1.0$, while the background is represented by a single NeRF network. Each NeRF expert is structured as a 7-layer MLP with 256 channels per layer. In training and inference, 256 coarse and 512 fine points are sampled per ray for the foreground network, whereas 128 coarse and 256 fine samples are sampled for the background network. The training utilizes 8 NVIDIA RTX 3090 GPUs, with a global batch size of 8192 rays in each iteration. The models are optimized by 500k iterations for each large-scale scene and then evaluated on validation images.
@@
In our Switch-NeRF++ framework, we utilize multi-resolution hash encoding with 16 levels, each level with 2 feature channels. The maximum hash map size is $2^{19}$. For the Mega-NeRF and Matrixcity datasets, we use the same hash experts. For the larger Block-NeRF dataset and UrbanBIS dataset, we validate our design of a heterogeneous mixture of hash experts (HMoHE) using both the same hash experts and heterogeneous hash experts. The hash gating network includes a multi-resolution hash encoding and a 3-layer MLP with 64 channels. We set 8 hash experts in our experiments for the foreground and one hash encoding for the background. The $\sigma$ prediction head is a 2-layer MLP and the color prediction head is a 3-layer MLP, both with 64 channels. We conduct our experiments on 4 NVIDIA RTX 3090 GPUs, with a batch size of 16384 rays per GPU. We use larger batch sizes to fully use the computation power of the GPUs since our network uses much less memory than MLP-based NeRF methods. We use coarse-to-fine sampling similar to Mega-NeRF, sampling 128 coarse samples and 128 fine samples along a ray in the foreground and half of them in the background.
@@
We use Adam optimizer to train Switch-NeRF and Switch-NeRF++ with a learning rate of $5\times 10^{-4}$ that exponentially decreases. The weight $\lambda$ of balance loss $\mathcal{L}_{b}$ is set as $5\times 10^{-4}$ for all experiments. We also use half-precision training and testing, similar to Mega-NeRF.
The detailed configuration and training pipeline of Switch-NeRF++ are shown in Algorithm 1. We configure the network with a hash-based gating module following the default Instant-NGP settings, and define $N=8$ hash experts with logarithmically spaced base and desired resolutions, enabling each expert to specialize in different scene scales. For each 3D point $\mathbf{x}$, it is processed by the gating network, one of the hash experts, and prediction heads to predict density and color. Volume rendering is applied along the ray to render the pixel color and compute $L_{2}$ loss. A gating balance loss is also used to balance the training.
@@
Baseline. We compare our method with the large-scale NeRF method Mega-NeRF [2]. We also adapt the Instant-NGP [5] into a large-scale setting, named INGP. We use a foreground and a background network in the baseline. Each of them consists of one multi-resolution hash encoding and prediction heads. We set the hash map size as $2^{22}$ in the foreground network so it has the same network capacity as our method. The sampling strategy is the same as our Switch-NeRF++.
@@
TABLE I: The benchmark performances of our Switch-NeRF++ on the Mega-NeRF [2] dataset. In this table, the INGP is the Instant-NGP [5] trained with 100k steps. Switch-NeRF++* is trained with 50k steps, and Switch-NeRF++ is trained with 100k steps. Our method can achieve state-of-the-art results with only about a training time of 5.5h, which is an 8x acceleration compared to Switch-NeRF [4] and a 5x acceleration compared to Mega-NeRF [2].
@@ 4.3 Overall Benchmark Results
Mega-NeRF dataset. We evaluate the accuracy and efficiency of our framework on Mega-NeRF [2] dataset. The accuracy and training time are reported in Table I. The results of Mega-NeRF is quoted from their papers. In Table I, Mega-NeRF and Switch-NeRF are trained with about 30h and 44h, respectively, for 500k steps. Instant-NGP (INGP) [5] is trained with about 10h for 100k steps. Switch-NeRF++* is our method trained with 5.5h for 50k steps. Switch-NeRF++ is our method trained with 11h for 100k steps. The results show that our MLP-based Switch-NeRF can clearly outperform its counterpart Mega-NeRF, validating the effectiveness of our learnable decomposition and Mixture of Experts. Switch-NeRF++* trained with only 5.5 hours can already achieve state-of-the-art results on almost all the metrics. This is an acceleration of $8\times$ compared to Switch-NeRF and $5\times$ compared to Mega-NeRF, proving the superior efficiency and accuracy of our method. Our Switch-NeRF++ clearly outperforms INGP with less training time, showing the great scalability of our Switch-NeRF++ framework. The training time of our method is comparable to INGP with the same steps. One reason is that we implement an efficient dispatching method by integrating point dispatching into the hash operation. Another reason is that we set a similar number of parameters in INGP as in our network. We provide qualitative results in Fig. 8. Our method can render images of higher quality with cleaner and sharper details. The main results on the Mega-NeRF dataset clearly demonstrate that our Switch-NeRF++ can achieve better accuracy with superior efficiency compared to Mega-NeRF, Switch-NeRF, and INGP.
UrbanBIS dataset. We evaluate the accuracy and scalability of our method on the very large UrbanBIS [10] dataset. Table II demonstrates that our heterogeneous hash experts (Heter.) significantly outperform all the compared methods. Notably, our heterogeneous version shows substantial improvements over our homogeneous version (Homo.) across all metrics (PSNR, SSIM, and LPIPS), emphasizing the importance of considering scene heterogeneity when scaling to larger scenes. These experiments strongly support the scalability of our design of HMoHE in Switch-NeRF++ when applied to very large-scale scenes. Fig. 9(a) illustrates that our method produces images with enhanced sharpness and details.
@@
4.4 Model Analysis
Effect of heterogeneous mixture of hash experts. We validate our heterogeneous hash design on Block-NeRF [3] and UrbanBIS [10] datasets. The scene in the Block-NeRF dataset covers a $1km$ street. The scene in the UrbanBIS dataset covers a $6.5km^{2}$ region. Both scenes are large enough to evaluate the scalability of our heterogeneous design. The results in Table II clearly demonstrate the superior performance of our heterogeneous design compared to our homogeneous version (Homo.), particularly on UrbanBIS. Each of the experts in Switch-NeRF++ (Homo.) uses the default setting of the encodings in Instant-NGP [5]. In contrast, Switch-NeRF++ (Heter.) incorporates experts with varying resolution ranges, with minimum resolutions of the experts exponentially increasing from 16 to 512 and maximum resolutions exponentially increasing from 2048 to 16384. The ablation results strongly emphasize the significance of considering scene heterogeneity when scaling to larger scenes and provide clear evidence for the scalability of our method. Fig. 9(b) visualizes the decomposition of the radiance field of UrbanBIS dataset of Switch-NeRF++ (Homo.). and Switch-NeRF++ (Heter.), and Fig. 10 visualizes the decomposition of the Mega-NeRF dataset. As shown in these figures, the point clouds of Switch-NeRF++ (Homo.) have similar or random sharpness, showing that their decomposition is not related to local scales. In contrast, the point clouds of Switch-NeRF++ (Heter.) demonstrate a pattern of increasing sharpness along with the increase of resolution. These visualizations further validate that our heterogeneous design can effectively capture the different distributions of the large-scale scene.
TABLE III: PNSR on MatrixCity [76] dataset. Switch-NeRF++ achieves an overall 1.75-point improvement in PSNR. The results of other methods are quoted from the MatrixCity paper. Please note that INGP also uses $2^{22}$ entries, whose number of parameters aligns with our network. These experiments demonstrate that our method can handle different types of scenes.
@@ 4.4 Model Analysis
To further analyze the heterogeneous representation learned by heterogeneous experts, we visualize the correlation between the density gradient distribution and the expert structure on the large UrbanBIS scene in Fig. 11. Our motivation for this analysis is that the density gradient at a 3D point reflects local density variation, indicating the local scene scale. The regions with less density variations can be represented by coarse-scale grids, while the regions with more density variations need fine-grained grids. In Fig. 11a, the distributions of gradients have no clear correlation with the experts. Different experts tend to cover a wide and homogeneous range of gradients (scales). However, in Fig. 11b of our heterogeneous model, we can clearly observe that the experts with lower resolutions tend to cover more regions of small gradients, while those with higher resolutions tend to cover more regions of large gradients. This effectively validates that our design can learn heterogeneous decomposition and representation of large-scale scenes.
@@
Figure 10: Visualization results of our decomposition of the radiance field on the Mega-NeRF datasets. Our heterogeneous designs can capture quite different distributions across different scenes, from coarser to finer structures, while our homogeneous designs handle similar distributions.
Gating network design in Switch-NeRF. We analyze different designs of the MLP-based gating network in Switch-NeRF [4]. Table VIII compares three configurations. The first is using a Linear gating, which processes the positional encoding PE(x) by a trainable linear layer. The second is removing the LayerNorm in the gating network. The third is training the network without the gating loss $L_{b}$. As shown in the table, Linear gating underperforms due to insufficient network capacity, validating that MLP+Norm design in our gating network can substantially improve reconstruction accuracy. Notably, removing the gating loss $L_{b}$ prevents the training from converging, highlighting that $L_{b}$ is necessary for stabilizing the optimization of our sparse MoE system. This finding aligned with prior works  [11, 64].
Full dispatch in Switch-NeRF. We compare the inference accuracy of using uniform dispatch or using our full dispatch proposed in Sec. 3.5.1. The Batch Prioritized Routing [65] and a capacity factor $C_{f}$ of $2.0$ or $4.0$ are used in the inference of the uniform dispatch. As shown in Table XII, uniform dispatch with $C_{f}=2.0$ suffers significant quality degradation. While increasing $C_{f}$ to $4.0$ partially improves the quality, it remains substantially inferior to full dispatch. Meanwhile, a larger $C_{f}$ introduces extra computational overhead due to zero padding. Visual comparisons in Fig. 6 further highlight that the uniform dispatch renders images with more artifacts, whereas our full dispatch produces complete and higher-quality outputs. This validates the effectiveness of full dispatch in inference of Switch-NeRF.
Figure 11: The visualization of the distribution of density gradient and the expert structure for the large UrbanBIS scene. In (a) Homo., the gradient distribution does not show a clear correlation with the experts. In contrast, in (b) Heter., when the expert resolution increases, the heterogeneous expert handles more points with larger gradients and fewer points with smaller gradients. This gradient analysis also aligns well with the expert point visualization, in which the point clouds in (b) Heter. demonstrate a pattern of increased sharpness along with the increase in expert resolution. This figure illustrates clearly that our heterogeneous structure can learn heterogeneous decomposition and representation.
TABLE VIII: Different gating network designs in Switch-NeRF on Building scene.
@@
TABLE IX: Comparison of Switch-NeRF and Mega-NeRF with different experts on accuracy, test time, test memory, and network parameter numbers. Switch-NeRF demonstrates better scalability than Mega-NeRF, increasing performance without a large increase in test memory and time.
@@
Effect of the hash-based gating network. We replace the hash-based gating network in Switch-NeRF++ with an MLP gating network similar to that in Switch-NeRF [4], which consists of 4 linear layers with a width of 256. The input of the MLP gating network is the positional encoding [1]. We train them with 100k steps. As shown in Table IV, the accuracy drops by a large margin with the MLP gating network. This may be because the MLP suffers from slow converging speed, and the representation power of the MLP gating network is inconsistent with the hash experts in Switch-NeRF. This experiment further confirms the importance of our design of the hash-based gating network.
TABLE X: The ablation results of Switch-NeRF using different Top-11, Top-22 and capacity factors. A larger $C_{f}$ or employing Top-22 operation in the gating network in training can improve the accuracy slightly while consuming more training time and memory.
@@
TABLE XI: Ablation results comparing Switch-NeRF++, Switch-NeRF [4], and Mega-NeRF [2] on memory usage and computation time during training and testing. Switch-NeRF++ achieves significantly faster training and testing speeds. Note that Switch-NeRF++ uses a 4x larger batch size than Switch-NeRF and Mega-NeRF. Therefore, it consumes slightly more training memory than Mega-NeRF but still less than Switch-NeRF. In testing, Switch-NeRF++ speed up rendering by 16x compared to Switch-NeRF [4] and by 13x compared to Mega-NeRF [2] with much less memory.
Method Test Train
$downarrow Mem.\downarrow Time\downarrow$
Mega-NeRF 6935M 87.9s 5124M 30.7h
Switch-NeRF 5847M 110s 10182M 42.5h
Switch-NeRF++ 2857M 6.65s 8707M 10.7h
TABLE XII: The testing time and accuracy of uniform dispatch with $C_{f}=2.0$ and $4.0$ and full dispatch. Full dispatch can achieve much better performance with less time usage.
@@
TABLE XIII: Impact of different numbers of gating operations on accuracy and efficiency in Switch-NeRF. Despite maintaining the same number of expert layers, Switch-NeRF with two gating operations fails to enhance rendering accuracy compared to a single gating operation, while requiring more memory usage and longer training and inference times.
@@
Figure 12: Comparison with large-scale Gaussian methods and Zip-NeRF. The zoomed-in image patches show that GS-based methods such as CityGaussian and Hier-GS tend to produce floating artifacts and blurred structures in challenging regions, while Switch-NeRF++ can produce high-fidelity and stable reconstructions.
Effect of the number of experts. We analyze the effect of the number of NeRF experts in both Switch-NeRF [4]. Table IX reveals that even using 4 experts, Switch-NeRF already matches the accuracy of Mega-NeRF with 8 experts. Scaling Switch-NeRF to 16 experts yields significant performance gains compared to 8 experts, while incurring minimal additional computational and memory overhead during inference. This contrasts with the Mega-NeRF with 16 experts, where Switch-NeRF not only scales much better and achieves a higher accuracy, but also reduces parameters and memory usage. The results demonstrate the superior scalability of Switch-NeRF in balancing model capacity and resource demands.
@@
TABLE XIV: Comparison with large-scale GS methods and Zip-NeRF. Experiment settings are aligned with VastGaussian.
Dataset Building Rubble Residence Sci-Art Campus
Metrics PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Hier-GS 22.24 0.711 0.217 24.97 0.715 0.226 23.27 0.775 0.175 26.60 0.857 0.155
Zip-NeRF 16.22 0.418 0.548 21.37 0.526 0.474 19.75 0.598 0.383 19.41 0.635 0.399 16.60 0.406 0.698
VastGaussian 23.50 0.804 0.130 26.92 0.823 0.132 24.25 0.852 0.124 26.81 0.885 0.121 26.00 0.816 0.151
CityGaussian 22.43 0.761 0.179 26.08 0.793 0.164 23.12 0.809 0.161 24.42 0.837 0.171
Switch-NeRF++ 22.41 0.633 0.290 26.02 0.680 0.281 24.00 0.730 0.224 27.40 0.823 0.190 24.37 0.594 0.429
Figure 13: The accuracy of our Switch-NeRF++ and other methods with respect to training steps. Our method continues to improve with more training steps. We also calculate the changing rates of expert assignment with respect to the training steps. In the early training, the changing rate is at a reasonably high value to enable the network to optimize the gate assignment. The changing rate decreases to about 0.05, showing that the gating network has converged. Note that the steps are x1k in the plot.
@@
TABLE XV: Additional ablation studies on the detailed configurations of Switch-NeRF++, evaluated using the PSNR metric (higher is better). Gumbel uses Gumbel-Softmax for expert selection. E16 denotes 16 experts; HS2 uses 8 experts with a doubled hash table size. E8,HS1/8 refers to 8 experts with 1/8 of the original hash table size, while INGP is an Instant-NGP baseline using a single hash grid with the same number of parameters as E8,HS1/8. HS1/2,F2 indicates half the hash table size with doubled feature dimensions. Head uses the same prediction head as Switch-NeRF. Ours shows the performance of our standard Switch-NeRF++ settings.
Gumbel E16 HS2 E8,HS1/8 INGP HS1/2,F2 Head Ours
18.22 23.55 23.35 22.32 21.41 23.22 23.31 23.25
Plane features without decomposition. The GP-NeRF [54] augments hash encodings with 2D ground plane features. It does not perform scene decomposition. The comparison results are shown in Table V. Switch-NeRF++ consistently outperforms GP-NeRF, especially on the very large UrbanBIS [10] dataset. These experiments validate the importance of our learnable heterogeneous decomposition on large-scale scenes to better exploit network parameters.
@@
Vanilla Gaussian splatting. We compare with the vanilla Gaussian splatting method (3D-GS) [44] in Table VII on large-scale datasets. The results demonstrate that the vanilla 3D-GS struggles on large-scale scenes. It relies on a point cloud for initialization, while point clouds in large-scale scenes are typically noisy, contain outliers, and have significantly varying densities. The original 3D-GS has no special designs for these problems, resulting in poor initialization and training. Our method incorporates the learnable heterogeneous decomposition specially designed for large-scale scenes and can get much better results. Moreover, our method does not need any point clouds for initialization, making it more suitable for various large-scale reconstruction tasks.
TABLE XVI: A detailed analysis of changes in memory usage and computational complexity with network designs in Switch-NeRF++. The training batch size and steps are aligned with the Switch-NeRF++ in our main results.
Hash Gate Effic. Disp. Hash Exp. & Head Train Test
Mem. Time Mem. Time
$\checkmark$ 18503M 27.4h 5461M 104.2s
$\checkmark$ $\checkmark$ 15859M 22.0h 4430M 71.4s
$\checkmark$ $\checkmark$ $\checkmark$ 8707M 10.7h 2857M 6.65s
Large-scale Gaussian splatting and Zip-NeRF. We compare with Zip-NeRF [80] and Large-scale Gaussian splatting methods such as Hierarchical-3DGaussians [56], CityGaussian [57], and Vast-Gaussian [59] in Table XIV. In these experiments, we align the training and evaluation settings of Switch-NeRF++ with Vast-Gaussian, such as the training and evaluation resolutions, the color correction in evaluation and the AlexNet version of LPIPS. We include results for Hierarchical-3DGaussians and Zip-NeRF using their open-source training and evaluation codes. We obtain the results of City-Gaussian using its released checkpoints. All evaluation metrics of these methods are also aligned with the Vast-Gaussian setting.
As shown in Table XIV, Switch-NeRF++ achieves significantly higher PSNR than Hierarchical-3DGaussians. Hierarchical-3DGaussians, which decomposes the scene via a tree structure and merges chunks by post-training, struggles to generalize across different scene types due to its rule-based design. Compared with CityGaussian, Switch-NeRF++ achieves comparable PSNR on Building and Rubble and performs notably better on Residence and Sci-Art (27.40 vs. 24.42). This suggests that rule-based large-scale Gaussian methods still face challenges in handling sparse supervision, noisy point clouds, and diverse scene types, which are common issues in large-scale settings. While Switch-NeRF++ underperforms Vast-Gaussian on some scenes (e.g., Building, Rubble, and Residence), it performs competitively and surpasses it on Sci-Art, which contains challenging, distant regions with sparse point clouds.
Although GS methods generally report better SSIM and LPIPS, this is partly because of their use of SSIM loss. This may also stem from the fundamental differences between GS-based rasterization and our NeRF-style ray-based rendering. GS-based methods are fundamentally point-based rendering based on sparse point clouds, which may focus more on sharp structures and may get higher SSIM scores. Our NeRF-style ray-based rendering may put more emphasis on global consistency and smoothness among 3D points. However, Fig. 12 shows that GS-based methods tend to produce visible artifacts in air or distant regions, which may be overlooked by SSIM. Our method can produce much cleaner and more globally consistent images.
While Zip-NeRF [80] also extends Instant-NGP for multi-scale rendering, it is not specifically designed for large-scale scenes. As shown in Table XIV and Fig. 12 (Paper Fig. 12), Zip-NeRF struggles in large-scale settings, whereas Switch-NeRF++ achieves superior results across various scenes.
Fig. 12 presents a qualitative comparison between Switch-NeRF++ and large-scale Gaussian methods. GS-based methods often suffer from artifacts in distant or low-density regions due to their reliance on high-quality input point clouds. In contrast, our method maintains robustness across these difficult areas.
These limitations of GS-based methods arise from their reliance on high-quality point clouds and dense image coverage. However, these conditions are rarely met in large-scale scene datasets, which typically contain sparse images considering the large areas they cover, and noisy and sparse point clouds. Additionally, the rule-based decomposition used in GS-based methods struggles to generalize across different scene types. As a result, large-scale GS methods still face challenges in these settings, whereas Switch-NeRF++ remains robust and competitive.
More ablation studies. Table XV presents additional ablations on the Residence scene. Using Gumbel-Softmax for expert selection leads to unstable training and poor results, confirming the effectiveness of the MoE structure. Increasing the number of experts (E16) yields greater accuracy gains than enlarging the hash map (HS2), likely due to improved specialization and better hash table utilization. In contrast, larger hash maps may result in under-utilized entries. Reducing the hash map size while doubling the feature dimension (“HS1/2,F2") to maintain parameter count shows minimal accuracy loss, indicating robustness to higher collision rates. Replacing the prediction head in Switch-NeRF++ with that of Switch-NeRF has little impact, suggesting that gains stem mainly from the heterogeneous hash expert design.